# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ H.1 Initial Results
With the information-theoretic tools from Section 3.1 in place, we proceed to state an initial non-asymptotic bound of [39] that we use as a starting point. Recalling that we already split the analysis into $\ell\leq\frac{k}{\log k}$ and $\ell>\frac{k}{\log k}$ (and here we consider the latter), we define $\ell_{\min}=1+\lfloor\frac{k}{\log k}\rfloor$ for notational convenience.
@@
$\mathbb{P}({\rm err})\leq\mathbb{P}\bigg{[}\bigcup_{(s_{\mathrm{dif}},s_{\mathrm{eq}})\,:\,|s_{\mathrm{dif}}|\geq\ell_{\min}}\bigg{\{}\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\leq\log{{p-k}\choose|s_{\mathrm{dif}}|}+\log\bigg{(}\frac{k}{\delta_{1}}{k\choose|s_{\mathrm{dif}}|}\bigg{)}\bigg{\}}\bigg{]}+\delta_{1}.$ (340)
@@
$\max_{\ell=\ell_{\min},\dotsc,k}\frac{\log{{p-k}\choose\ell}+\log\big{(}\frac{k}{\delta_{1}}{k\choose\ell}\big{)}}{I_{\ell}^{n}(1-\delta_{2})}\leq 1,$ (341)
@@
$\mathbb{P}\big{[}\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\leq(1-\delta_{2})I_{\ell}^{n}\big{]}\leq\psi_{\ell}(n,\delta_{2}),$ (342)
@@
$\mathbb{P}({\rm err})\leq\sum_{\ell=\ell_{\min}}^{k}{k\choose\ell}\psi_{\ell}(n,\delta_{2})+\delta_{1}.$ (343)
@@
We can take $\delta_{1}\to 0$ sufficiently slowly to have an asymptotically negligible impact in (341), so the key step in ensuring $\mathbb{P}({\rm err})\to 0$ is showing that the summation in (343) approaches 0. We will take $\delta_{2}$ to be a constant arbitrarily close to zero, as was done in [39], [^8] For low $\ell$ values it should instead be taken as a parameter to be optimized [39], but we are only using these techniques for the $-\ell$ regime. and we will show that the condition (341) simplifies to a requirement on $n$ corresponding to the first term in (4).
Naturally, to simplify (341) we need to characterize the behavior of $I_{\ell}^{n}$. For the Bernoulli design, the following asymptotic characterization was given in [39].
Lemma 9.
@@
$I_{\ell}^{n}=\left(\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log\frac{1-\rho}{\rho}\right)\big{(}1+o(1)\big{)}.$ (344)
@@
$I_{\ell}^{n}=ne^{-(1-\alpha)\nu}\big{(}H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho)\big{)}\big{(}1+o(1)\big{)}.$ (345)
We will show that the same result holds for the near-constant weight design in Lemma 10 below.
@@ H.2 Bernoulli Design
The following concentration term was shown to be valid (i.e., (342) holds) in [39, Prop. 3]:
$\psi_{\ell}(n,\delta_{2})=2\exp\bigg{(}-\frac{n(\delta_{2}I_{\ell})^{2}}{4(8+\delta_{2}I_{\ell})}\bigg{)},$ (346)
where $I_{\ell}=I_{\ell}^{n}/n$ is the mutual information associated with a single test. Under this choice, it is shown in [39, Prop. 7(ii)] that $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$ as long as $n=\Omega(k\log\frac{p}{k})$ (regardless of the implied constant). Moreover, in [39, Sec. C.2] it is shown for $\nu=\log 2$ that the maximum in (341) is asymptotically attained by $\ell=k$, and that the limiting behavior exactly corresponds to the first term in (4).
A slight additional requirement in our work is that we need to handle general $\nu>0$, rather than only $\nu=\log 2$, with the non-trivial step being to show that the maximum in (341) is still asymptotically attained by $\ell=k$ [^9] The proof of $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$ is not impacted by changing $\nu$, since doing so only amounts to changing constant factors in the exponent in (346). Fortunately, this follows from another well-known result in the literature [32, Thm. 3] (see also [45, Lemma 2]) that we now proceed to outline.
First note that for $\ell>\frac{k}{\log k}$ and $k=o(p)$, the second term in the numerator of (341) is asymptotically negligible compared to the first. We further have $\log{p-k\choose\ell}=\big{(}\ell\log\frac{p}{\ell}\big{)}(1+o(1))=\big{(}\ell\log\frac{p}{k}\big{)}(1+o(1))$, so the maximization over $\ell$ simplifies to maximizing $\frac{\ell}{I_{\ell}^{n}}$, or equivalently minimizing $\frac{I_{\ell}^{n}}{\ell}$. The above-mentioned results in [32, Thm. 3] and [45, Lemma 2] directly state that $\ell=k$ minimizes $\frac{I_{\ell}^{n}}{\ell}$ (for a broader class of noise models that includes ours), as desired.
Having established that $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$ and that $\ell=k$ asymptotically attains the maximum in (341), the desired threshold on $n$ follows by writing the numerator therein as $\big{(}k\log\frac{p}{k}\big{)}(1+o(1))$ and the denominator as $n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))\times(1-\delta_{2})$ (see (345) with $\alpha=1$), and recalling that $\delta_{2}$ can be arbitrarily small. This establishes the first term in (4).
@@ H.3 Near-Constant Weight Design
For the near-constant weight design, we again adopt the choices of $\delta_{1}$ and $\delta_{2}$ described above for the Bernoulli design. Notice that the study of the condition (341) has no direct relation to the choice of $\psi_{\ell}$. Thus, to simplify (341), we can apply the exact same analysis as the Bernoulli case as long as we can show that the mutual information terms $I_{\ell}^{n}$ have the same asymptotic behavior; this is done in the following lemma, whose proof is given in Appendix J.2.
Lemma 10.
@@
$I_{\ell}^{n}=\left(\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log\frac{1-\rho}{\rho}\right)\big{(}1+o(1)\big{)}.$ (347)
@@
$I_{\ell}^{n}=ne^{-(1-\alpha)\nu}\big{(}H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho)\big{)}\big{(}1+o(1)\big{)}.$ (348)
From this lemma and the above discussion, we deduce that (341) again leads to the first term in (16). It only remains to show that $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$.
While we cannot directly use (346), the proof of [39, Prop. 7(ii)] reveals that we do not precisely need (346), but rather, the term inside the exponent can be scaled be any finite constant factor. Since $I_{\ell}=O(1)$ and $\delta_{2}=\Theta(1)$, the denominator in (346) is dominated by the constant term (to within a constant factor), and thus, what we seek is a tail bound of the form $e^{-\Theta(nI_{\ell}^{2})}=e^{-\Theta((I_{\ell}^{n})^{2}/n)}$. We provide this in the following lemma, whose proof is given in Appendix J.1.
Lemma 11.
(Concentration Bound for $\frac{\ell}{k}\to\alpha$, Near-Constant Weight Design) For the noisy group testing problem under near-constant design, consider sequences $k\to\infty$ and $\ell\geq 1$ (indexed by $p$) such that $\frac{\ell}{k}\to\alpha\in[0,1]$. For any given $\delta$ smaller than a quantity depending on $(\alpha,\nu,\rho)$, the following holds for sufficiently large $p$ and any $s_{\mathrm{dif}}\subset S$ with cardinality $\ell$:
$\mathbb{P}\Big{(}\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})\leq(1-\delta)I_{\ell}^{n}\Big{)}\leq\exp\Big{(}-\frac{C_{\rho,\nu,\alpha}(I_{\ell}^{n})^{2}\delta^{2}}{n}\Big{)},$ (349)
where $C_{\rho,\mu,\alpha}>0$ is a constant depending on $(\rho,\nu,\alpha)$, and $I_{\ell}^{n}$ satisfies (347) when $\alpha=0$, or (348) when $\alpha\in(0,1]$.
The proof is given in Appendix J.1. With this result in place, the rest of the analysis again follows that of the Bernoulli design to establish the first term in (11).
@@ Appendix I $-\ell$ Converse Analysis for Both Designs
Recall that the thresholds in (4) and (11) contain two terms; this appendix concerns only the first, which corresponds to the $-\ell$ regime.
@@ I.1 Bernoulli Design
For the Bernoulli design, the first term in (4) was already given in [39, Thm. 7 & App. C], and mainly comes down to studying the lower tail probability of the information density $\imath^{n}(\mathbf{X}_{s},\mathbf{Y})$ (i.e., (18) with $s_{\mathrm{eq}}=\emptyset$, corresponding to $\ell=k$). Here establishing concentration is relatively straightforward due of the i.i.d. nature of the Bernoulli design. We omit any further details, since (i) they are already known from [39], and (ii) for the near-constant weight design (analyzed below) we will follow a similar structure and we will give more detail for it.
@@ I.2 Near-Constant Weight Design
Unlike the Bernoulli design, the first term in (11) does not readily follow from [39], since their analysis was specific to the Bernoulli design. However, it turns out that we can get the desired result by combining certain tools from [39] with a fairly simple analysis of the information density.
@@
$\mathbb{P}({\rm err})\geq\mathbb{P}\bigg{(}\imath^{n}(\mathbf{X}_{s},\mathbf{Y})\leq\log\Big{(}\delta_{1}{p\choose k}\Big{)}\bigg{)}-\delta_{1}$ (350)
for any $\delta_{1}>0$, where $s=[k]$ (without loss of generality), and $\imath^{n}(\mathbf{X}_{s},\mathbf{Y})=\log\frac{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})}{\mathbb{P}(\mathbf{Y})}$ is a shorthand for the information density (340) with $s_{\mathrm{eq}}=\emptyset$ (and $s_{\mathrm{dif}}=[k]$). Hence, its mean is $I_{k}^{n}$ (i.e., $I_{\ell}^{n}$ from (19) with $\ell=k$), which behaves as $n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))$ by Lemma 10.
@@
$\mathbb{P}\big{(}\imath^{n}(\mathbf{X}_{s},\mathbf{Y})\leq(1+\delta_{2})I_{k}^{n}\big{)}\to 1.$ (351)
To do so, write $\imath^{n}(\mathbf{X}_{s},\mathbf{Y})=\log\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})-\log\mathbb{P}(\mathbf{Y})$ and handle the two terms separately. For the first term, we note that given $\mathbf{X}_{s}$, the sequence $\mathbf{Y}$ has independent entries depending on the ${\rm Bernoulli}(\rho)$ noise, from which a simple concentration argument (e.g., Hoeffding’s inequality) gives $\log\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})=-nH_{2}(\rho)(1+o(1))$ with probability approaching one. For the second term, we decompose $\mathbb{P}(\mathbf{Y})=\mathbb{P}(V)\mathbb{P}(\mathbf{Y}|V)$, where $V$ is the number of 1s in $\mathbf{Y}$. We will utilize two useful lemmas regarding $V$, the first of which is the following concentration bound.
Lemma 12.
@@
See (363) in the proof of Lemma 11 below (Appendix J). Note that (363) considers general $(s_{\mathrm{eq}},s_{\mathrm{dif}})$ and defines $V$ to be the number of 1s in tests that contain no items from $s_{\mathrm{eq}}$, but this reduces to the above definition when $s_{\mathrm{eq}}=\emptyset$ (and thus $\alpha=1$ in (363)). ∎
@@
Combining the preceding two findings with the fact that $I_{k}^{n}=n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))$ (Lemma 10), we see that in order to establish (351), it only remains to show that the impact of the remaining term $\log P(V)$ is asymptotically negligible; namely, it should scale as $o(n)$ to be insignificant compared to the leading $\Theta(n)$ terms. This is stated in the following lemma, which is proved in Appendix J.
Lemma 13.
@@
Intuitively, this result holds because $V$ is produced via binomial random variables (due to the independent noise variables); it can readily be deduced from anti-concentration bounds (Lemma 1) that for $U\sim{\rm Bin}(N,q)$ with $N\to\infty$ and constant $q\in(0,1)$, we have that $P(U=u)\geq e^{-o(N)}$ whenever $u=Nq(1+o(1))$ (which holds with high probability). Since the full details are slightly more technical without being much more insightful than this intuition, we defer the formal proof of Lemma 13 to Appendix J.3.
Combining the above findings, we have established that (351) holds for arbitrarily small $\delta_{2}$. Moreover, in (350), we can take $\delta_{1}$ to decay with $p$ sufficiently slowly such that $\log\big{(}\delta_{1}{p\choose k}\big{)}=\big{(}k\log\frac{p}{k}\big{)}(1+o(1))$ (i.e., the impact of $\delta_{1}$ is asymptotically negligible), and comparing this $\big{(}k\log\frac{p}{k}\big{)}(1+o(1))$ term with the term $I_{k}^{n}=n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))$ in (351), we deduce the desired threshold on $n$ given by the first term in (11).

### 23 Oct 2024 v2

@@ H.1 Initial Results
With the information-theoretic tools from Section 3.1 in place (recalling in particular the information density $\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})$ and mutual information $I_{\ell}^{n}$ with $\ell=|s_{\mathrm{dif}}|$), we proceed to state an initial non-asymptotic bound from [39] that we use as a starting point. Recalling that we already split the analysis into $\ell\leq\frac{k}{\log k}$ and $\ell>\frac{k}{\log k}$ (and here we consider the latter), we define $\ell_{\min}=1+\lfloor\frac{k}{\log k}\rfloor$ for notational convenience.
@@
$\mathbb{P}({\rm err})\leq\mathbb{P}\bigg{[}\bigcup_{(s_{\mathrm{dif}},s_{\mathrm{eq}})\,:\,|s_{\mathrm{dif}}|\geq\ell_{\min}}\bigg{\{}\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\leq\log{{p-k}\choose|s_{\mathrm{dif}}|}+\log\bigg{(}\frac{k}{\delta_{1}}{k\choose|s_{\mathrm{dif}}|}\bigg{)}\bigg{\}}\bigg{]}+\delta_{1}.$ (H.1)
@@
$\max_{\ell=\ell_{\min},\dotsc,k}\frac{\log{{p-k}\choose\ell}+\log\big{(}\frac{k}{\delta_{1}}{k\choose\ell}\big{)}}{I_{\ell}^{n}(1-\delta_{2})}\leq 1,$ (H.2)
@@
$\mathbb{P}\big{[}\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\leq(1-\delta_{2})I_{\ell}^{n}\big{]}\leq\psi_{\ell}(n,\delta_{2}),$ (H.3)
@@
$\mathbb{P}({\rm err})\leq\sum_{\ell=\ell_{\min}}^{k}{k\choose\ell}\psi_{\ell}(n,\delta_{2})+\delta_{1}.$ (H.4)
@@
We can take $\delta_{1}\to 0$ sufficiently slowly to have an asymptotically negligible impact in (H.2), so the key step in ensuring $\mathbb{P}({\rm err})\to 0$ is showing that the summation in (H.4) approaches 0. We will take $\delta_{2}$ to be a constant arbitrarily close to zero, as was done in [39], [^9] For low $\ell$ values it should instead be taken as a parameter to be optimized [39], but we are only using these techniques for the $-\ell$ regime. and we will show that the condition (H.2) simplifies to a requirement on $n$ corresponding to the first term in (1.4).
Naturally, to simplify (H.2) we need to characterize the behavior of $I_{\ell}^{n}$. For the Bernoulli design, the following asymptotic characterization was given in [39].
Lemma H.1.
@@
$I_{\ell}^{n}=\left(\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log\frac{1-\rho}{\rho}\right)\big{(}1+o(1)\big{)}.$ (H.5)
@@
$I_{\ell}^{n}=ne^{-(1-\alpha)\nu}\big{(}H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho)\big{)}\big{(}1+o(1)\big{)}.$ (H.6)
We will show that the same result holds for the near-constant weight design in Lemma H.2 below.
@@ H.2 Bernoulli Design
The following concentration term was shown to be valid (i.e., (H.3) holds) in [39, Prop. 3]:
$\psi_{\ell}(n,\delta_{2})=2\exp\bigg{(}-\frac{n(\delta_{2}I_{\ell})^{2}}{4(8+\delta_{2}I_{\ell})}\bigg{)},$ (H.7)
where $I_{\ell}=I_{\ell}^{n}/n$ is the mutual information associated with a single test. Under this choice, it is shown in [39, Prop. 7(ii)] that $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$ as long as $n=\Omega(k\log\frac{p}{k})$ (regardless of the implied constant). Moreover, in [39, Sec. C.2] it is shown for $\nu=\log 2$ that the maximum in (H.2) is asymptotically attained by $\ell=k$, and that the limiting behavior exactly corresponds to the first term in (1.4).
A slight difference in our work is that we need to handle general $\nu>0$, rather than only $\nu=\log 2$, with the non-trivial step being to show that the maximum in (H.2) is still asymptotically attained by $\ell=k$ [^10] The proof of $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$ is not impacted by changing $\nu$, since doing so only amounts to changing constant factors in the exponent in (H.7). Fortunately, this follows from another well-known result in the literature [32, Thm. 3] (see also [45, Lemma 2]) that we now proceed to outline.
First note that for $\ell>\frac{k}{\log k}$ and $k=o(p)$, the second term in the numerator of (H.2) is asymptotically negligible compared to the first. We further have $\log{p-k\choose\ell}=\big{(}\ell\log\frac{p}{\ell}\big{)}(1+o(1))=\big{(}\ell\log\frac{p}{k}\big{)}(1+o(1))$, so the maximization over $\ell$ simplifies to maximizing $\frac{\ell}{I_{\ell}^{n}}$, or equivalently minimizing $\frac{I_{\ell}^{n}}{\ell}$. The above-mentioned results in [32, Thm. 3] and [45, Lemma 2] directly state that $\ell=k$ minimizes $\frac{I_{\ell}^{n}}{\ell}$ (for a broader class of noise models that includes ours), as desired.
Having established that $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$ and that $\ell=k$ asymptotically attains the maximum in (H.2), the desired threshold on $n$ follows by writing the numerator therein as $\big{(}k\log\frac{p}{k}\big{)}(1+o(1))$ and the denominator as $n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))\times(1-\delta_{2})$ (see (H.6) with $\alpha=1$), and recalling that $\delta_{2}$ can be arbitrarily small. This establishes the first term in (1.4).
@@ H.3 Near-Constant Weight Design
For the near-constant weight design, we adopt the same choices of $\delta_{1}$ and $\delta_{2}$ as those for the Bernoulli design. Notice that the study of the condition (H.2) has no direct relation to the choice of $\psi_{\ell}$. Thus, to simplify (H.2), we can apply the exact same analysis as the Bernoulli case as long as we can show that the mutual information terms $I_{\ell}^{n}$ have the same asymptotic behavior; this is done in the following lemma, whose proof is given in Appendix J.2.
Lemma H.2.
@@
$I_{\ell}^{n}=\left(\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log\frac{1-\rho}{\rho}\right)\big{(}1+o(1)\big{)}.$ (H.8)
@@
$I_{\ell}^{n}=ne^{-(1-\alpha)\nu}\big{(}H_{2}(e^{-\alpha\nu}\star\rho)-H_{2}(\rho)\big{)}\big{(}1+o(1)\big{)}.$ (H.9)
From this lemma and the above discussion, we deduce that (H.2) again leads to the first term in (1.15). It only remains to show that $\sum_{\ell>\frac{k}{\log k}}{k\choose\ell}\psi_{\ell}(n,\delta_{2})\to 0$.
While we cannot directly use (H.7), the proof of [39, Prop. 7(ii)] reveals that we do not precisely need (H.7), but rather, the term inside the exponent can be scaled be any finite constant factor. Since $I_{\ell}=O(1)$ and $\delta_{2}=\Theta(1)$, the denominator in (H.7) is dominated by the constant term (to within a constant factor), and thus, what we seek is a tail bound of the form $e^{-\Theta(nI_{\ell}^{2})}=e^{-\Theta((I_{\ell}^{n})^{2}/n)}$. We provide this in the following lemma, whose proof is given in Appendix J.1.
Lemma H.3.
(Concentration Bound for $\frac{\ell}{k}\to\alpha$, Near-Constant Weight Design) For the noisy group testing problem under near-constant design, consider sequences $k\to\infty$ and $\ell\geq 1$ (indexed by $p$) such that $\frac{\ell}{k}\to\alpha\in[0,1]$. For any given $\delta$ smaller than a quantity depending on $(\alpha,\nu,\rho)$, the following holds for sufficiently large $p$ and any $(s_{\mathrm{dif}},s_{\mathrm{eq}})$ with $|s_{\mathrm{dif}}|=\ell$:
$\mathbb{P}\Big{(}\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})\leq(1-\delta)I_{\ell}^{n}\Big{)}\leq\exp\Big{(}-\frac{C_{\rho,\nu,\alpha}(I_{\ell}^{n})^{2}\delta^{2}}{n}\Big{)},$ (H.10)
where $C_{\rho,\mu,\alpha}>0$ is a constant depending on $(\rho,\nu,\alpha)$, and $I_{\ell}^{n}$ satisfies (H.8) when $\alpha=0$, or (H.9) when $\alpha\in(0,1]$.
The proof is given in Appendix J.1. With this result in place, the rest of the analysis again follows that of the Bernoulli design to establish the first term in (1.11).
@@ Appendix I $-\ell$ Converse Analysis for Both Designs
Recall that the thresholds in (1.4) and (1.11) contain two terms; this appendix concerns only the first, which corresponds to the $-\ell$ regime.
@@ I.1 Bernoulli Design
For the Bernoulli design, the first term in (1.4) was already given in [39, Thm. 7 & App. C], and mainly comes down to studying the lower tail probability of the information density $\imath^{n}(\mathbf{X}_{s},\mathbf{Y})$ (i.e., (3.1) with $s_{\mathrm{eq}}=\emptyset$, corresponding to $\ell=k$). Here establishing concentration is relatively straightforward due of the i.i.d. nature of the Bernoulli design. We omit any further details, since (i) they are already known from [39], and (ii) for the near-constant weight design (analyzed below) we will follow a similar structure and we will give more detail for it.
@@ I.2 Near-Constant Weight Design
Unlike the Bernoulli design, the first term in (1.11) does not readily follow from [39], since their analysis was specific to the Bernoulli design. However, it turns out that we can get the desired result by combining certain tools from [39] with a fairly simple analysis of the information density.
@@
$\mathbb{P}({\rm err})\geq\mathbb{P}\bigg{(}\imath^{n}(\mathbf{X}_{s},\mathbf{Y})\leq\log\Big{(}\delta_{1}{p\choose k}\Big{)}\bigg{)}-\delta_{1}$ (I.1)
for any $\delta_{1}>0$, where we implicitly condition on the defective set being $S=s=[k]$ (without loss of generality), and $\imath^{n}(\mathbf{X}_{s},\mathbf{Y})=\log\frac{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})}{\mathbb{P}(\mathbf{Y})}$ is a shorthand for the information density (H.1) with $s_{\mathrm{eq}}=\emptyset$ (and $s_{\mathrm{dif}}=[k]$). Hence, its mean is $I_{k}^{n}$ (i.e., $I_{\ell}^{n}$ from (3.2) with $\ell=k$), which behaves as $n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))$ by Lemma H.2.
@@
$\mathbb{P}\big{(}\imath^{n}(\mathbf{X}_{s},\mathbf{Y})\leq(1+\delta_{2})I_{k}^{n}\big{)}\to 1.$ (I.2)
To do so, we write $\imath^{n}(\mathbf{X}_{s},\mathbf{Y})=\log\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})-\log\mathbb{P}(\mathbf{Y})$ and handle the two terms separately. For the first term, we note that given $\mathbf{X}_{s}$, the sequence $\mathbf{Y}$ has independent entries depending on the ${\rm Bernoulli}(\rho)$ noise, from which a simple concentration argument (e.g., Hoeffding’s inequality) gives $\log\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})=-nH_{2}(\rho)(1+o(1))$ with probability approaching one. For the second term, we decompose $\mathbb{P}(\mathbf{Y})=\mathbb{P}(V)\mathbb{P}(\mathbf{Y}|V)$, where $V$ is the number of 1s in $\mathbf{Y}$. We will utilize two useful lemmas regarding $V$, the first of which is the following concentration bound.
Lemma I.1.
@@
See (J.12) in the proof of Lemma H.3 in Appendix J. Note that (J.12) considers general $(s_{\mathrm{eq}},s_{\mathrm{dif}})$ and defines $V$ to be the number of 1s in tests that contain no items from $s_{\mathrm{eq}}$, but this reduces to the above definition when $s_{\mathrm{eq}}=\emptyset$ (and thus $\alpha=1$ in (J.12)). ∎
@@
Combining the preceding two findings with the fact that $I_{k}^{n}=n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))$ (Lemma H.2), we see that in order to establish (I.2), it only remains to show that the impact of the remaining term $\log P(V)$ is asymptotically negligible; namely, it should scale as $o(n)$ to be insignificant compared to the leading $\Theta(n)$ terms. This is stated in the following lemma, which is proved in Appendix J.
Lemma I.2.
@@
Intuitively, this result holds because $V$ is produced via binomial random variables (due to the independent noise variables); it can readily be deduced from anti-concentration bounds (Lemma A.1) that for $U\sim{\rm Bin}(N,q)$ with $N\to\infty$ and constant $q\in(0,1)$, we have that $P(U=u)\geq e^{-o(N)}$ whenever $u=Nq(1+o(1))$ (which holds with high probability). Since the full details are slightly more technical without being much more insightful than this intuition, we defer the formal proof of Lemma I.2 to Appendix J.3.
Combining the above findings, we have established that (I.2) holds for arbitrarily small $\delta_{2}$. Moreover, in (I.1), we can take $\delta_{1}$ to decay with $p$ sufficiently slowly such that $\log\big{(}\delta_{1}{p\choose k}\big{)}=\big{(}k\log\frac{p}{k}\big{)}(1+o(1))$ (i.e., the impact of $\delta_{1}$ is asymptotically negligible), and comparing this $\big{(}k\log\frac{p}{k}\big{)}(1+o(1))$ term with the term $I_{k}^{n}=n(H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho))(1+o(1))$ in (I.2), we deduce the desired threshold on $n$ given by the first term in (1.11).